OpenStudy (anonymous):

Find a formula for the derivative of m(x)=(1/(x+1)) using the difference quotient.

zepdrix (zepdrix):

The `difference quotient` is given by (I'll write it using the function m so it's more applicable to us):\[\Large \frac{m(x+h)-m(x)}{h}\]

zepdrix (zepdrix):

So ummmm

zepdrix (zepdrix):

Are you familiar with the `Limit Definition` of the derivative?\[\Large \lim_{h\to0}\frac{f(x+h)-f(x)}{h}\]Learned about that yet? :) That's what we'll need to use here.

OpenStudy (anonymous):

Ya, to be honest I really just am having a problem with the algebra and fractions involved. I did it out and got : \[f'(x)= -(x^2+2x+hx+h+1)\]

zepdrix (zepdrix):

Ya the algebra is a bit tricky, let's see what went wrong.

zepdrix (zepdrix):

\[\Large f(x)=\frac{1}{1+x}, \qquad\qquad f(x+h)=\frac{1}{1+x+h}\]So we'll need to plug these into our formula.

zepdrix (zepdrix):

\[\Large f'(x)\quad=\quad \lim_{h\to0}\frac{\frac{1}{1+x+h}-\frac{1}{1+x}}{h}\]

zepdrix (zepdrix):

Setup look ok so far? :)

OpenStudy (anonymous):

Yes the next thing I did was in the numerator mulitply by (1+x+h/1+x+h) and (1+x/1+x) so they had the same denominator.

zepdrix (zepdrix):

\[\Large f'(x)\quad=\quad \lim_{h\to0}\frac{\left(\color{royalblue}{\dfrac{(1+x)-(1+x+h)}{(1+x+h)(1+x)}}\right)}{h}\]

zepdrix (zepdrix):

Hmm something like that I guess :o

zepdrix (zepdrix):

So it's easier to read, we can move the denominator down to the basement with the h.\[\Large f'(x)\quad=\quad \lim_{h\to0}\frac{(1+x)-(1+x+h)}{h(1+x+h)(1+x)}\]

zepdrix (zepdrix):

Hopefully that little trick wasn't too confusing :D

OpenStudy (anonymous):

That's what I did, I have a hard time simplifying it from there. I'm glad I did it right so far though.

zepdrix (zepdrix):

In the numerator remember to distribute the negative to each term in the second set of brackets. We'll get some nice cancellations up there.\[\Large f'(x)\quad=\quad \lim_{h\to0}\frac{\cancel{1+x}\cancel{-1-x}-h}{h(1+x+h)(1+x)}\]

OpenStudy (anonymous):

Oh ok, I see it now! Thanks so much!

zepdrix (zepdrix):

cool :)